What is the sphericity of a hollow cylinder whose length is 10 times of its own diameter?


A) 0.14


B) 0.21


C) 0.37           D) 0.69


E) None of the above


1
Expert's answer
2021-08-20T01:37:03-0400

"\phi_s= \frac{6 V_P}{D_pS_p}\\
V_p=\frac{\pi d^2 h}{4} =10*\frac{\pi d^3 }{4}\\"

Sp can be found by

"\frac{\pi}{6}D_p^3=\frac{\pi}{4}d_p^3\\
D_p=6d^3\\
D_p= 1.14d_p\\
S_p= 1-D_p\\
S_p= 1-1.14\\
S_p=0.14"